OpenStudy (anonymous):

I toss a coin and rolls a number cube 100 times. How many times should I expect to have the coin show heads and roll a 1 or a 2?

OpenStudy (kropot72):

The probability of a head is 1/2. The probability of a 1 or a 2 is 1/3. Therefore the probability of a head and a 1 or a 2 is given by: \[\frac{1}{2}\times\frac{1}{3}=you\ can\ calculate\] When you have found the fraction for the probability, you can calculate the expected number of times out of 100 tosses/rolls that heads and 1 or 2 are obtained by multiplying 100 by the probability. Hint: You will need to round the result to an integer.

OpenStudy (kropot72):

@justonestudent Are you there?

OpenStudy (anonymous):

@kropot72 I got about 17

OpenStudy (kropot72):

Good work! You are correct.

OpenStudy (anonymous):

Thanks so much!

OpenStudy (kropot72):

You're welcome:)